Hello,
Steve traveled 600 miles at a certain speed. Had he gone 20mph faster, the trip would have taken 1 hour less. Find the speed of his vehicle.
Thank you

Answer by checkley77(12569)

D=RT OR T=D/R
T=600/R
(T-1)=600/(R+20)
T=600/(R+20)+1
600/R=600/(R+20)+1
600/R=(600+R+20)/R+20
600(R+20)=R(R+620)
600R+12000=R^2+620R
R^2+620R-600R-12000=0
R^2+20R-1,2000=0
(R-100)(R+120)=0
R-100=0
R=100 MPH ANS. FOR THE LONGEST TIME TRIP
R=100+20=120 MPH ANS. FOR THE 1 HOUR SHORTENED TRIP.
600/100=6 hrs.
600/120=5 hrs.